OpenStudy (anonymous):

At 1 P.M., ship A is 150 km west of ship B. Ship A is sailing east at 35 km/h and ship B is sailing north at 25 km/h. How fast is the distance between the ships changing at 5 P.M.? (Round your answer to one decimal place.

OpenStudy (amistre64):

how fast is the distance changing...thats me calculus ;)

OpenStudy (amistre64):

there is a 4hour difference, so our trianlge has to reflect that fact..

OpenStudy (anonymous):

i keep getting 53.7

OpenStudy (anonymous):

but its wrong

OpenStudy (amistre64):

x^2 + y^2 = r^2 Dx 2x + Dy 2y = Dr 2r ; dicide the 2s Dx x + Dy y = Dr r Dr = Dx x + Dy y //r Dx =35; Dy =25. lets find x, y and r :)

OpenStudy (amistre64):

y = 25(4) x = 35(4) +150

OpenStudy (anonymous):

y did u miltiply by 4?

OpenStudy (anonymous):

ooooo i c

OpenStudy (amistre64):

r = sqrt(100^2 + 190^2)

OpenStudy (amistre64):

4 hours right :)

OpenStudy (anonymous):

i multiplied them by 5

OpenStudy (anonymous):

ok dont tell me answer..i wanna c if i can get this

OpenStudy (amistre64):

Dr = 35(100) + 25(190) ----------------- sqrt(10000 + 190^2)............you might to close your eyes for this then :)

OpenStudy (anonymous):

where did the 190 come from?

OpenStudy (anonymous):

the 100 came from multipying the 25 by 100 right

OpenStudy (amistre64):

150 + 4(35) = 150 + 140 = ...290.....

OpenStudy (anonymous):

i mean 25 by 4

OpenStudy (amistre64):

25(4) yes = 100

OpenStudy (anonymous):

oo ok

OpenStudy (amistre64):

see, its good to check my work; after all, im just an idiot an disguise ;)

OpenStudy (anonymous):

so is it 49.7

OpenStudy (amistre64):

dunno, I stopped the math cause you didnt wanna know lol

OpenStudy (amistre64):

Dr = 35(290) + 25(100) ----------------- looks better to me eh sqrt(10000 + 290^2)

OpenStudy (amistre64):

10150 +2500 7650 ------------ = ---------- = 24.91... is what I get sqrt(94100) sqrt(94100)

OpenStudy (amistre64):

i had to adjust for my own stupidity :)

OpenStudy (anonymous):

it said it was wrong :(

OpenStudy (amistre64):

..... well the Dr part is correct, double check the distances to be sure.

OpenStudy (anonymous):

for last step i did it differently, i did (140+100)/290 *(35+25)

OpenStudy (amistre64):

Dr = Dw w + Dn n //r lets call it north and west...

OpenStudy (amistre64):

or e for east.... im going blind i tell ya..BLIND!!

OpenStudy (amistre64):

Dr = De e + Dn n //r De = 35; Dn = 25 e = 150 + 4(35); n = 4(25) e = 290; n = 100 Dr = 35(290) + 25(100) //r r = sqrt(290^2 + 100^2) = sqrt(84100 + 10000) = sqrt(94100) Dr = 10150 + 2500 12650 ------------- = ------- sqrt(94100) sqrt(94100) Dr = 41.2378

OpenStudy (amistre64):

that should be it :)

OpenStudy (anonymous):

ok thank u!!

OpenStudy (amistre64):

just hope its right :)

OpenStudy (anonymous):

u really are a life saver!!

OpenStudy (amistre64):

thanx ;)